Aggregator

Cryo Daily Update

2 days 5 hours ago
Log Number: 4551718Logbook(s): Body: 

• CHL Control Room receptacle relocation- completed
• CHL Control Room cabinet removal - completed
• HDR Cold Box maintenance - in progress
• CHL2 vertical section of 4K Cold Box’s diffusion pump electrical maintenance - in progress
• CHL2’s 4K Cold Box (both sections) – complete mechanical maintenance on vacuum pumps
• CHL2’s 2K Cold Box started purging through process piping to the recovery system and installed a relief valve

jwilson

C100-02R Acceptance Testing Swing Shift

2 days 6 hours ago
Log Number: 4551703Logbook(s): Body: 

Completed Qo measurements for cavities 2 and 4. All Qo's are complete.
Cavity 2 Qo at 8 MV/m = 1.4E10, Qo at Emaxop = 5.8E9
Cavity 4 Qo at 8 MV/m = 1.2E10, Qo at Emaxop = 9.2E9
Pressure Sensitivity measured for both cavities ~ 220 Hz / torr

drury

MOLLER install 11/3/2026

2 days 7 hours ago
Log Number: 4551696Logbook(s): Tags: Entry Makers: jonesBody: 

• Starting welding of the aluminum in the MD cable support stands.
• Started cutting the material for the DAQ cable trays
• Completed the shielding for the beam dump camera. Picture attached
• Target group completed the vacuum test of the h2 piping through the truck ramp
• Target group completed the leak test of the new cryo transfer line.

jones

COMMENT ON PSS General maintenance tasks for SAM 2026

2 days 8 hours ago
Log Number: 4551693Logbook(s): Body: 

Completed junction box wiring and cable labeling for the emergency switches in the North Access Room. Verified that the switches are functioning properly.

Installed buzzers for the Tornado Shelter emergency switches at Doors 1 and 2 in the South Access Room. Added new cables and switches to provide power. Junction box wiring and cable labeling remain to be completed.

L. Krylach, S. Bruhwel

krylach

COMMENT ON Coordination of CEBAF Wien filter Jobs

2 days 8 hours ago
Log Number: 4551692Logbook(s): Body: 

Today's progress updates:

1. Mamun submitted ATLIS (#121909) to baseline high voltage operation prior to magnet disassembly
2. DC power plans disconnects of H-Wien and V-Wien magnets, likely this week, will forecast at 8am mtg (no conflict with #1)
3. CIS and I&C will walk electrode high voltage system, attach WCD's (disconnects happen after #1)

grames

COMMENT ON Coordination of CEBAF Wien filter Jobs

2 days 8 hours ago
Log Number: 4551691Logbook(s): Body: 

A small group met today to discuss modifying the CEBAF Wien filters, to add air cooling holes to the yokes.

It was a great meeting, discussed scope, hazards, resources, roles and responsibilities.

Here's our path forward, with the lead person responsible to coordinate their WCD's.

1. Joe G., will coordinate with DC (Onish) and I&C (Keith) to isolate and remove stored energy hazards for magnet repair and modifications, clear the area for repair jobs, after a magnet is reassembled coordinate restoring power for testing

2.

grames

Hall C daily summary

2 days 8 hours ago
Log Number: 4551690Logbook(s): Body: 

Today Hall C technicians worked on.

Painting PCS skids.
Machining MOLER Barite shield wall assembly parts.
Fabricating concrete anchor pieces.
Welding on ENGY mobile detector stand.
Printing and test-fitting and modifying NUCAT parts
Assembling filament extrusion project.
Group training for EPAS V5
Gathered appropriate seals for downstream beam pipe.
Prepped beam pipe for assembly.
No cross division work planned for 11/04/2026

jnines

COMMENT ON Install 8 kW circulators in 1L09

2 days 9 hours ago
Log Number: 4551686Logbook(s): Body: 

Succeeded getting 1L09 to 14kV after clearing issues with door interlocks, HPA IOC restore, and high body current (KBCU) trips.

Cavities #6 and #7 issues with high KBCU trips seemed were caused by connectors needing reterminaion on Klystron connection board. At least for cavity #7, faulty connection was reterminated. Will attempt to replace connector for cavity 6 on next opportunnity. Cavity 6 is bypassed until getting a spare connector.

Will continue testing on Wednesday.

xgomez

COMMENT ON Laser Room Rack Upgrade

2 days 10 hours ago
Log Number: 4551680Logbook(s): Body: 

We pulled back all cables under the rack, most just go out the existing floor level pass through. There were many cables just coiled up. All cables that were removed were stacked on pile to be surveyed. We could not get any cable to push back through the pass through. In the meantime, we will try to ID cables that are ready to remove after Shukui looks them over (or a list, not many have any label).

The rack is wide open and ready for the BNC and fiber pull (lower open area on front); and the Ethernet switch install on the rear anywhere high.

scovert

iocsoftnl08rf restored

2 days 10 hours ago
Log Number: 4551675Logbook(s): Tags: Body: restorelib 3.0.1 Target restore date was 2026-11-03 13:56:51 The output from the myRestore execution was as follows: iocsoftnl08rf <NOTE> rf3int: Provided macro *IOC* is never used <NOTE> rf3int: Provided macro *REBOOTPATH* is never used <NOTE> rf3hpa: Provided macro *IOC* is never used <NOTE> rf3hpa: Provided macro *REBOOTPATH* is never used <NOTE> rf3stp: Provided macro *IOC* is never used <NOTE> rf3stp: Provided macro *REBOOTPATH* is never used <NOTE> rf3fcc1: Provided macro *IOC* is never used <NOTE> rf3f
alarms

iocsoftnl08rf restored

2 days 11 hours ago
Log Number: 4551672Logbook(s): Tags: Body: restorelib 3.0.1 Target restore date was 2026-11-03 13:41:22 The output from the myRestore execution was as follows: iocsoftnl08rf <NOTE> rf3int: Provided macro *IOC* is never used <NOTE> rf3int: Provided macro *REBOOTPATH* is never used <NOTE> rf3hpa: Provided macro *IOC* is never used <NOTE> rf3hpa: Provided macro *REBOOTPATH* is never used <NOTE> rf3stp: Provided macro *IOC* is never used <NOTE> rf3stp: Provided macro *REBOOTPATH* is never used <NOTE> rf3fcc1: Provided macro *IOC* is never used <NOTE> rf3f
alarms

Shift Information

Shift Day
CEBAF Program
MCC Crew Chief Richardson
MCC Operators
Program Deputy Merz
Announcements

Access Status

Recent PSS History

recent pss history graph